Determine the probability of drawing an even-numbered card or a card with a red suit from a standard deck of cards. Leave your answer in a reduced fraction.


1
Expert's answer
2022-02-14T16:10:27-0500

There are 26 red cards, 5*4=20 even-numbered cards and 5*2=10 even-numbered red cards.

"P=\frac{26+20-10}{52}=\frac{36}{52}=\frac{9}{13}" .